# HAIGEN: Towards Human-AI Collaboration for Facilitating Creativity and Style Generation in Fashion Design

### v1

@@ Abstract.
The process of fashion design usually involves sketching, refining, and coloring, with designers drawing inspiration from various images to fuel their creative endeavors. However, conventional image search methods often yield irrelevant results, impeding the design process. Moreover, creating and coloring sketches can be time-consuming and demanding, acting as a bottleneck in the design workflow. In this work, we introduce HAIGEN (Human-AI Collaboration for GENeration), an efficient fashion design system for Human-AI collaboration developed to aid designers. Specifically, HAIGEN consists of four modules. T2IM, located in the cloud, generates reference inspiration images directly from text prompts. With three other modules situated locally, the I2SM batch generates the image material library into a certain designer-style sketch material library. The SRM recommends similar sketches in the generated library to designers for further refinement, and the STM colors the refined sketch according to the styles of inspiration images. Through our system, any designer can perform local personalized fine-tuning and leverage the powerful generation capabilities of large models in the cloud, streamlining the entire design development process. Given that our approach integrates both cloud and local model deployment schemes, it effectively safeguards design privacy by avoiding the need to upload personalized data from local designers. We validated the effectiveness of each module through extensive qualitative and quantitative experiments. User surveys also confirmed that HAIGEN offers significant advantages in design efficiency, positioning it as a new generation of aid-tool for designers.
@@ 3.2. System Overview
3.3. Privacy Protection
As depicted in Table 1, existing generation methods like Midjourney (Midjourney, 2022) and DALL-E (Ramesh et al., 2021) lack the concept of cloud-local integration. They rely entirely on cloud server resources for computations. Users must upload their data to these servers, where the corresponding content is generated based on the input and then downloaded locally for viewing. This poses a significant privacy risk, as designers are required to upload their private works and inspirations to these servers. DataMix (Liu et al., 2020), while utilizing the cloud-local collaboration framework and leveraging cloud computing resources to alleviate local computing burdens. However, it necessitates the uploading of local data models to the cloud. Despite the introduction of a privacy protection solution, potential privacy concerns persist.
Table 1. The comparison of several different methods for privacy protection.
Midjourney (Midjourney, 2022) DALL-E (Ramesh et al., 2021) DataMix (Liu et al., 2020) HAIGEN (Ours)
Computation (on Cloud) 100% 100% 87% 72%
Computation (on Local) 0% 0% 13% 28%
[…]
Design Availability Internet Internet Internet LAN
In contrast, our proposed cloud-local collaboration framework strategically allocates tasks based on complexity and privacy considerations. Models requiring extensive computational resources and posing no privacy concerns, such as inspiration image generation models, are deployed in the cloud. Conversely, models related to designer privacy and requiring minimal computational resources are deployed locally. This approach harnesses the potent generation capabilities of large cloud models while eliminating the privacy risks associated with local data. Moreover, concerning Design Availability, outputs generated by techniques like Midjourney (Midjourney, 2022) are showcased in a public chat window, enabling others to view the generated results. In contrast, results produced by our approach are exclusively accessible to the respective designer or shared within the Local Area Network (LAN) of the design studio. This discrepancy highlights a crucial distinction in privacy protection between our method and others. Notably, through our framework, designers only need to download inspiration images generated in the cloud, without uploading any image-related private data. Thus, it restricts local data upload, further safeguarding designer privacy.
@@ 4.1. Stable Diffusion Model-based Text-to-Image Cloud Module
In our Text-to-Image Cloud Module (T2IM), we generate a variety of high-quality clothing renderings based on text prompts, offering designers an intuitive sense of the designed clothing’s quality. It possesses the following characteristics: (i) Generates clothing try-on renderings directly from detailed text descriptions and allows for the adjustment of parameters such as character background, image style, and size, etc. (ii) Requires only a small amount of dataset and computing resources for fine-tuning the model with other style images. (iii) Can generate mixed-style image effects by combining multiple style models. As shown in Fig. 5, our T2IM primarily comprises two components: one is a model based on Stable Diffusion (SD) model (Rombach et al., 2022), and the other involves fine-tuning the SD model using LoRA (Hu et al., 2021) and ControlNet (Zhang et al., 2023a). This fine-tuning accelerates the model training process and enhances the detailed information of the generated images.
@@ 4.2. Capture Personalized Designer-Style Image-to-Sketch Local Module
In previous works (Canny, 1986; Kim et al., 2020; Liu et al., 2021b), the task of image-to-sketch generation can be summarized as extracting contour and detail features from the input image. However, the sketch generation model we proposed assists designers in quickly completing sketches with minimal computational resource consumption. It ensures that the generated sketches maintain their unique artistic style by learning from the designer’s historical hand-drawn works. This is because each designer has their own unique artistic style and conveys their perception of important elements in the sketches from various perspectives through their personal drawing techniques. To illustrate this, we present the framework of our Image-to-Sketch Local Module (I2SM) in Fig. 6, which primarily consists of three components. These components include the standard pre-trained VGG-16 (Simonyan and Zisserman, 2014) for extracting multi-level features, the Adaptive Personalized Style Normalization (APSN) Module for capturing and blending different designer styles, and the DownSampling Multi-Feature Fusion (DSMFF) Module in Decoder for multi-level feature reconstruction.
@@ 4.3. Cloud Image-based Local Personalized Sketch Recommendation Module
In this module, our objective is to swiftly identify sketches from the local sketch material library that closely match the cloud image. This process assists designers in refining the recommended sketch templates, ultimately facilitating the creation of a sketch-based image style suitable for coloring. Illustrated in Fig. 7, we present the workflow of the personalized sketch recommendation system. It involves obtaining a substantial collection of personalized sketch material libraries through a locally pre-trained sketch-to-image generation model and image material library (Sec. 4.2). We generate numerous images of the clothing model try-on in various scenes through a cloud-based pre-trained text-to-image generation model and the input reference text prompt (Sec. 4.1). Then, we retrieve the corresponding clothing image and recommend the most similar sketch materials from the personalized sketch material library based on the clothing image. Designers can further refine the recommended sketch materials and input them into our Style Transfer Module (Sec. 4.4).
@@ 4.4. Style Transfer Module based on Local Personalized Sketch and Cloud Image
At this stage, our proposed Style Transfer Module (STM) merges the sketch obtained locally refined with the image acquired from the cloud model, generating a customized sketch image imbued with the stylistic elements of the cloud image, as present in Fig. 8. Our STM is built around the core component of the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020), which governs the model’s diffusion generation process by converting sketch and image features into explicit and implicit supervision. Additionally, we enhance the conditional supervision of the diffusion process through our proposed Channel Cross Attention Module (CCAM), ensuring the production of high-quality, realistic images. This approach empowers designers to swiftly apply personalized colors to their sketches, facilitating the selection of the most suitable color style for their work, thereby enhancing designer efficiency. By utilizing private, locally stored sketches, we ensure the preservation of designers’ creative concepts, effectively safeguarding their privacy.
@@ 5.6.1. Qualitative Results
Fig. 12 illustrates the impressive performance of HAIGEN in recommending sketch templates. The first column represents the input reference image, followed by ten columns showing the top-10 recommended sketch templates from our built sketch library, most closely resembling the reference image. Notably, the recommended sketches closely resemble the real sketches, appearing prominently in the lineup. Furthermore, other sketch templates also exhibit striking similarities with the reference images. For instance, similarities in clothing styles can be observed in the second and third rows, and details such as the neckline of the clothes and pockets in the first row are replicated in the recommended templates. Although the first-ranked recommended sketch template does not precisely match the reference image, the resemblance between this recommendation and both the reference image and the real sketch template is evident. These results convincingly highlight the exceptional performance of our Sketch Recommendation Module in recommending sketch templates.
@@ 5.7.1. Quantitative Results
Table 3 illustrates the quantitative evaluation of our method using five metrics to compare the performance with other approaches. While our method may not exhibit a substantial advantage in model inference speed, this can be attributed to the inherent slow generation speed caused by the principle of the diffusion-based model itself. It is worth noting that our method exhibits clear advantages across all indicators, particularly excelling in LPIPS and FID, aligning with human eye perception and reaching new state-of-the-art levels. Specifically, our HAIGEN employs the DDIM algorithm with 100 samples, demonstrating substantial advantages. While the results with 20 and 50 samples exhibit slightly reduced accuracy, the benefits still stand out. For the 200 samples result, the increased time cost for interfacing doesn’t provide a proportional improvement in image quality, making it less cost-effective. Furthermore, the comparison indicates that our proposed CCAM enhances the overall performance in all metrics. This improvement is attributed to its nature as an attention module, utilizing cross-attention for enhanced feature fusion between feature channels. The matrix operations in the other branch of CCAM facilitate simple aggregation of feature details. These quantitative results effectively demonstrate the effectiveness of our approach in sketch-to-image style transfer.
@@
Methods Inference HAIFashion
(items/s) PSNR $\uparrow$ SSIM $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$
Baselines AdaIN (Huang and Belongie, 2017) 46.1538 17.5611 0.3168 0.0302 0.1016 67.3568
Self-Sup (Liu et al., 2021b) 16.2162 15.0941 0.3787 0.0446 0.1732 61.4071
StyleMe (Wu et al., 2023) 42.8571 22.8712 0.4576 0.0099 0.0651 30.7081
UCAST (Zhang et al., 2023b) 20.0000 18.8464 0.1144 0.0230 0.1275 43.2199
DiffusionRig (Ding et al., 2023) 2.1583 15.0146 0.1857 0.0359 0.1119 40.1118
Ablation w/o CCAM 0.4399 22.1428 0.3625 0.0076 0.0363 19.8801
20 samples 1.9355 23.1550 0.2425 0.0058 0.0591 30.9403
50 samples 0.7916 22.9912 0.2748 0.0060 0.0350 19.9451
200 samples 0.1926 23.9765 0.5000 $\textbf{0.0048}_{7}$ $0.0334_{4}$ 18.1717
HAIGEN (Ours) 0.4000 23.9430 0.3973 $0.0049_{3}$ $\textbf{0.0333}_{7}$ 18.1367
@@ 6.2. Survey Results
7. Discussion
7.1. Industry Impact
In our earlier user studies (Sec. 3.1), we discovered that designers now prefer using AI tools like Midjourney (Midjourney, 2022) to enhance their design process, significantly improving their design efficiency. These AI tools enable designers to generate images based on detailed text descriptions, facilitating the effective expression of their creative ideas. Moreover, contributing to the integration of multiple modules, our HAIGEN system offers comprehensive functionality for the entire design process, benefiting both novice and professional designers alike. Designers can further enhance works generated by the HAIGEN system through secondary development using other design software or platforms, thereby achieving superior design effects. Despite concerns regarding the emergence of AI, such as copyright and originality issues, and the potential for designers to become overly dependent on AI, we believe that judicious use of AI can offer numerous benefits to designers. Therefore, our focus in this paper is on leveraging AI as a tool to augment human creativity and design processes, fostering collaborative creation between humans and AI. While AI aids in enhancing efficiency and providing new ideas and inspiration, the ultimate decision-making and creative power remains in the hands of designers. Specifically, HAIGEN provides designers with inspiration references (Sec. 4.1) or material templates (Sec. 4.2) and enables quick previews of coloring effects (Sec. 4.4) for their works, etc. Throughout these processes, designers play an active role in drawing and refining works based on AI-generated outputs. This is achieved through a strategic combination of deployment segregation and collaborative capabilities between HAIGEN’s cloud large models and local small models. Such an approach effectively addresses concerns regarding design copyright infringement and leakage, while also mitigating the risk of excessive dependence on AI.
7.2. Limitation and Future Work
Although our system provides significant assistance to designers throughout the whole design process, there are still some weaknesses that need to be improved. Firstly, our system cannot currently mimic human-like thinking in generating sketches, instead relying on direct formation from the outline of images. Secondly, it cannot generate sketches of varying complexity tailored to the preferences of different designers. One potential solution for the first issue is to employ Bezier curves to simulate the drawing process of each stroke in the sketch, thereby achieving a more hand-drawn appearance. However, this approach introduces new challenges, particularly in the coloring process. Generating concrete images from abstract sketches poses a significant challenge as there is no strict one-to-one correspondence between the strokes of the sketch generated using Bezier curves and the corresponding image contours. This remains a problem that current algorithms struggle to solve effectively. As for the second issue, a feasible solution is to introduce a separate hyperparameter to control the complexity of generated sketches. This would allow designers to tailor the output to meet their specific design needs and facilitate targeted refinement. In summary, our future research will focus on exploring more general and applicable AI-assisted design solutions to address these challenges and further enhance the capabilities of our system.

### v3

@@ Abstract.
The process of fashion design usually involves sketching, refining, and coloring, with designers drawing inspiration from various images to fuel their creative endeavors. However, conventional image search methods often yield irrelevant results, impeding the design process. Moreover, creating and coloring sketches can be time-consuming and demanding, acting as a bottleneck in the design workflow. In this work, we introduce HAIGEN (Human-AI Collaboration for GENeration), an efficient fashion design system for Human-AI collaboration developed to aid designers. Specifically, HAIGEN consists of four modules. T2IM, located in the cloud, generates reference inspiration images directly from text prompts. With three other modules situated locally, the I2SM batch generates the image material library into a certain designer-style sketch material library. The SRM recommends similar sketches in the generated library to designers for further refinement, and the STM colors the refined sketch according to the styles of inspiration images. Through our system, any designer can perform local personalized fine-tuning and leverage the powerful generation capabilities of large models in the cloud, streamlining the entire design development process. Given that our approach integrates both cloud and local model deployment schemes, it effectively safeguards design privacy by avoiding the need to upload personalized data from local designers. We validated the effectiveness of each module through extensive qualitative and quantitative experiments. User surveys also confirmed that HAIGEN offers significant advantages in design efficiency, positioning it as a new generation of aid-tool for designers.  [^1] Accepted by Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (ACM IMWUT/UbiComp 2024). Please refer to the formal version for more details: https://dl.acm.org/doi/10.1145/3678518.
@@ 4.1. Stable Diffusion Model-based Text-to-Image Cloud Module
In our Text-to-Image Cloud Module (T2IM), we generate a variety of high-quality clothing renderings based on text prompts, offering designers an intuitive sense of the designed clothing’s quality. As shown in Fig. 5, our T2IM primarily comprises two components: one is a model based on Stable Diffusion (SD) model (Rombach et al., 2022), and the other involves fine-tuning the SD model using LoRA (Hu et al., 2021) and ControlNet (Zhang et al., 2023a). This fine-tuning accelerates the model training process and enhances the detailed information of the generated images.
@@ 4.2. Capture Personalized Designer-Style Image-to-Sketch Local Module
In previous works (Canny, 1986; Kim et al., 2020; Liu et al., 2021b), the task of image-to-sketch generation can be summarized as extracting contour and detail features from the input image. However, our proposed sketch generation model aims to assist designers in quickly completing hand-drawn sketch works, ensuring that the generated sketches possess the unique artistic style of the designer. This is because each designer has their own unique artistic style and conveys their perception of important elements in the sketches from various perspectives through their personal drawing techniques. To illustrate this, we present the framework of our Image-to-Sketch Local Module (I2SM) in Fig. 6, which primarily consists of three components. These components include the standard pre-trained VGG-16 (Simonyan and Zisserman, 2014) for extracting multi-level features, the Adaptive Personalized Style Normalization (APSN) Module for capturing and blending different designer styles, and the DownSampling Multi-Feature Fusion (DSMFF) Module in Decoder for multi-level feature reconstruction.
@@ 4.3. Cloud Image-based Local Personalized Sketch Recommendation Module
In this module, our goal is to identify sketches within the local sketch material library that closely match the input cloud image, thereby further improving the existing materials and finally realizing sketch-based image style transfer for coloring. Illustrated in Fig. 7, we present the workflow of the personalized sketch recommendation system. It involves obtaining a substantial collection of personalized sketch material libraries through a locally pre-trained sketch-to-image generation model and image material library (Sec. 4.2). We generate numerous images of the outfit worn in various scenes through a cloud-based pre-trained text-to-image generation model and the input reference text prompt (Sec. 4.1). Then, we retrieve the corresponding clothing image and recommend the most similar sketch materials from the personalized sketch material library based on the clothing image. Designers can further refine the recommended sketch materials and input them into our Style Transfer Module (Sec. 4.4).
@@ 4.4. Style Transfer Module based on Local Personalized Sketch and Cloud Image
At this stage, our proposed Style Transfer Module (STM) merges the sketch obtained locally refined with the image acquired from the cloud model, generating a customized sketch image imbued with the stylistic elements of the cloud image, as present in Fig. 8. This approach empowers designers to quickly apply personalized color to their sketches, facilitating the selection of the most fitting color style for their work and thereby enhancing designer efficiency. The use of a private, locally located sketch ensures that the designer’s creative ideas remain secure, effectively safeguarding their privacy. Our STM is built around the core component of the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020), with the generation of DDIM guided by the addition of sketches and images as content and style conditions. Furthermore, we enhance the conditions for the diffusion process through our proposed Channel Cross Attention Module (CCAM), ensuring the production of high-quality, realistic images.
@@ 5.6.1. Qualitative Results
Fig. 12 illustrates the impressive performance of HAIGEN in recommending sketch templates. The first column represents the input reference image, followed by ten columns showing the top-10 recommended sketch templates from our built sketch library, most closely resembling the reference image. Notably, the recommended sketches closely resemble the real sketches, appearing prominently in the lineup. Furthermore, other sketch templates also exhibit striking similarities with the reference images. For instance, similarities in clothing styles can be observed in the second and third rows, and details such as the neckline of the clothes and pockets in the first row are replicated in the recommended templates. Although the first-ranked recommended sketch template does not precisely match the reference image, the resemblance between this recommendation and both the reference image and the real sketch template is evident. These results convincingly highlight the exceptional performance of our Sketch Recommendation Module in recommending sketch templates
@@ 5.7.1. Quantitative Results
Table 2 illustrates the quantitative evaluation of our method using five metrics to compare the performance with other approaches. It is worth noting that our method exhibits clear advantages across all indicators, particularly excelling in LPIPS and FID, aligning with human eye perception and reaching new state-of-the-art levels. Specifically, our HAIGEN employs the DDIM algorithm with 100 samples, demonstrating substantial advantages. While the results with 20 and 50 samples exhibit slightly reduced accuracy, the benefits still stand out. For the 200 samples result, the increased time cost for interfacing doesn’t provide a proportional improvement in image quality, making it less cost-effective. Furthermore, the comparison indicates that our proposed CCAM enhances the overall performance in all metrics. This improvement is attributed to its nature as an attention module, utilizing cross-attention for enhanced feature fusion between feature channels. The matrix operations in the other branch of CCAM facilitate simple aggregation of feature details. These quantitative results effectively demonstrate the effectiveness of our approach in sketch-to-image style transfer.
@@
Methods HAIFashion
PSNR $\uparrow$ SSIM $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$
Baselines AdaIN (Huang and Belongie, 2017) 17.5611 0.3168 0.0302 0.1016 67.3568
Self-Sup (Liu et al., 2021b) 15.0941 0.3787 0.0446 0.1732 61.4071
StyleMe (Wu et al., 2023) 22.8712 0.4576 0.0099 0.0651 30.7081
UCAST (Zhang et al., 2023b) 18.8464 0.1144 0.0230 0.1275 43.2199
DiffusionRig (Ding et al., 2023) 15.0146 0.1857 0.0359 0.1119 40.1118
Ablation w/o CCAM 22.1428 0.3625 0.0076 0.0363 19.8801
20 samples 23.1550 0.2425 0.0058 0.0591 30.9403
50 samples 22.9912 0.2748 0.0060 0.0350 19.9451
200 samples 23.9765 0.5000 $\textbf{0.0048}_{7}$ $0.0334_{4}$ 18.1717
HAIGEN (Ours) 23.9430 0.3973 $0.0049_{3}$ $\textbf{0.0333}_{7}$ 18.1367